# MetaTool Benchmark: Deciding Whether to Use Tools and Which to Use

### 18 Jan 2024 v4

@@ 1 Introduction
Tool-empowered large language models (LLMs) (Qin et al., 2023a; b; Patil et al., 2023; Ruan et al., 2023; Liang et al., 2023) have recently attracted widespread attention. An important milestone for LLMs marching toward intelligent agents (Park et al., 2023; Liu et al., 2023b; Zhu et al., 2023) is the flexible use of tools (e.g., APIs (Qin et al., 2023b; Rapid, 2023) and plugins (OpenAI, 2023d)) to fulfill users’ requirements. By utilizing tools, LLMs can obtain real-time data, such as getting the latest weather forecast (GPTStore, 2023); enhance interactions with users, like helping users book flight tickets (Deng et al., 2023); and better deal with uncertain questions by querying knowledge bases (Li et al., 2023c; Hu et al., 2023) or Internet (Lazaridou et al., 2022). The capacity to use tools enables LLMs to break through their own limitations, acquire external information, and thereby make more accurate and effective responses, providing users with better service.
Previous research has focused on how to enhance the ability of LLMs to use tools, including training models with instruction related to tool usage (Qin et al., 2023b; Tang et al., 2023; Patil et al., 2023; Schick et al., 2023), or augmenting the model’s problem-solving capabilities for domain-specific tasks through external APIs (Yang et al., 2023c). A typical process of employing LLMs to use tools is illustrated in Figure 1. Initially, users input a question (i.e., query) that triggers the tool usage. Based on prior research (Yang et al., 2023c; Qin et al., 2023b), under the ReAct (Yao et al., 2022) prompt approach, the process of using tools can be divided into four stages: Firstly, LLMs consider whether to employ a tool (①) and if so, which tools to select (②). The tool selection process involves directly having LLMs choose from a provided tool list (Yang et al., 2023c; Ruan et al., 2023) or selecting via a retriever (Qin et al., 2023b). Next, LLMs configure the users’ input as tool parameters (③), then handle the results from the tool (④), and finally return the outcomes to the user.
@@ 2.1 Preliminary & Required Abilities
In this section, we first introduced the composition of the ToolE dataset, outlining how we generated user queries related to tools. Subsequently, we explained how we set up the evaluation tasks, including tool usage awareness and tool selection. The evaluation mainly requires LLMs with the following properties and abilities: (1) Less hallucination and sycophancy. The awareness of tool usage can reflect the truthfulness about whether an LLM has a clear understanding of its own capabilities (e.g., realizing its capability limitation about what problems it cannot solve well and using tools for assistance), thereby helping to mitigate issues of hallucination (Ji et al., 2023) and sycophancy (Wei et al., 2023). (2) Recommendation and retrieval. Moreover, existing research has tentatively explored the potential of LLMs in applications like LLM-based recommendation systems (e.g., tool recommendation for users) (Gao et al., 2023; Wang et al., 2023e; Dai et al., 2023). In LLM-as-agent scenarios, LLMs usually need to select the specific tool according to the text description (Park et al., 2023; Shen et al., 2023; Ruan et al., 2023), actually is a kind of information retrieval (Sun et al., 2023), making the ability of tool selection crucial. (3) Task-level abilities. In MetaTool, we set four tasks as shown in Table 12. Incorporating similar tools for selection (i.e., Task 1) requires a high-level semantic comprehension for LLMs, and tool selection in specific scenarios tests the flexibility of LLMs when using tools in different scenarios (e.g., finance (Wu et al., 2023a) and biomedical domain (Zhang et al., 2023; Wang et al., 2023c)). Task 3 aims to explore the internal hallucination and reliability extent of LLMs when using tools and Task 4 is designed to evaluate the inference ability (e.g., order of using multiple tools) (Creswell et al., 2022) of LLMs.
@@ 2.2.1 Dataset Generation
Multi-tool queries generation. Unlike single-tool queries, we generate multi-tool queries only after addressing the overlapped issue because it is challenging to map the original labels to new labels in the case of multi-label (i.e., multi-tool) situations. Here, we only consider queries related to two tools. We observe that if we obtain combinations of two tools by iterating through all the tools (i.e., C iterations, where is the size of the toolset), there would be many tool combinations that are not practical (i.e., rarely encountered in daily life, such as the combination of fortune-telling tools and currency exchange tools). Therefore, we select the top 15 most popular tools from the toolset, and for each pair of tools, we generate 5 queries. We determine the popularity of a tool based on the number of tools it is merged with, as shown in Appendix A.5. The multi-tool queries we generate can be divided into two types: The first category pertains to situations where tools are employed in parallel, indicating that the utilization of each tool operates independently of the others. The second category deals with cases where tools are used causally, signifying that the deployment of one tool may be contingent upon the outcomes of a preceding tool. Detailed prompt templates can be found in Appendix D.2. Similar to single-tool queries, we also manually verified multi-tool queries to ensure the combination of tools was reasonable and the query of the tool corresponded to the tool description.
@@ 2.3.1 Awareness of Tool Usage
In this part (i.e., Thought (①)), we aim to investigate the awareness of tool usage of LLMs; that is, whether LLMs can resort to external tools when they encounter problems they cannot solve. To this end, we need to construct the test set with both positive and negative samples. Positive samples are the queries that can not be solved by LLMs themselves and need tool usage, whereas negative samples are queries that can be directly solved by LLMs and therefore do not necessitate tool usage. For positive samples, we selected a subset of samples from ToolE and conducted manual validation to confirm whether they would trigger LLMs to use the tool (the process of which is detailed in the Appendix B). As for negative samples, we select three recent instruction datasets, including instructions about downstream tasks (Wang et al., 2022), common-sense questions (Talmor et al., 2019), and high-quality instructions used in LIMA (Zhou et al., 2023). Similarly, we conducted manual verification to ensure that these requests can be resolved by LLMs’ intrinsic capabilities. Specifically, we use the prompt with a query to inquire the LLMs whether need to employ a tool or not, and the output of LLMs should be either "yes" or "no".
@@ 2.3.2 Tool Selection
Sub-task 2: tool selection in specific scenarios. The objective of this task is to simulate how LLMs perform using tools when they act as controllers of a system (Shen et al., 2023) faced with different scenarios. As LLMs are widely applied across various domains like biomedical domain (Zhang et al., 2023) and educational domain (Kasneci et al., 2023), in scenarios where the system caters to diverse demographics or professions (e.g., software engineer (Qian et al., 2023)), its set of tools also varies. This task allows us to explore the performance disparities of LLMs in selecting different kinds of tools, essentially highlighting a form of bias inherent to LLMs (Ferrara, 2023). In such cases, this task examines how effectively LLMs utilize the tools. Given a query with its label , we specify the tool list containing candidates according to its corresponding scenarios. This task consists of two types of scenarios: the first one is the popularity-related scenario, and the second one is the group-related scenario. For the popularity-related scenario, we have selected the 5, 10, and 15 most popular tools based on the number of tools it’s merged with (refer to Table 9 in the Appendix A.5 for details.) to construct the tool list. As for the group-related scenario, we chose six usual occupations or identities and manually curated a tool list consisting of the 10 most relevant tools for each of them (see Table 10 in the Appendix A.5 for details).
Sub-task 3: tool selection with possible reliability issues. The reliability of LLMs’ tool selection is of utmost importance. However, issues like hallucination (Ji et al., 2023) and sycophancy (Wei et al., 2023) within LLMs’ responses will negatively impact their selection of tools. Therefore, we introduce sub-task 3. In this task, given a query and its corresponding tool , we need to construct the tool list and ensure . This aims to assess whether LLMs can answer questions honestly and avoid issues like choosing non-existent tools or selecting unrelated tools. It should be noted that this task closely resembles real-world scenarios, as not all existing tools capable of addressing user queries are present in the tool list controlled by LLMs. To be specific, we obtain the embedding of ’s description and get the top-k similar tools about as the way in Task 1. Then we randomly sample tools from the rest tool set to construct , denoted as where and . Overall, we remove the ground-truth tool of query and s most similar tools to keep the tools in not related to as much as possible.
@@ A.3 Guidelines for human validation
Calculation-related tools. ToolE contained numerous tools related to calculations. For simple calculations (e.g., ‘What is the value of sin 30 degrees?’ or ‘7 * 9 = ?’), LLMs can perform them without the need for a tool. However, for complex calculations, recent research (Yuan et al., 2023) suggests that LLMs still perform poorly. For queries corresponding to calculation-related tools, we removed queries involving simple calculations and retained those involving complex calculations.
@@ A.4 Flexibility of ToolE
Sometimes, when an LLM undergoes specialized training in a particular domain (Wu et al., 2023b; Wang et al., 2023a; Qi et al., 2023; Wang et al., 2023b; Yang et al., 2023a), its capabilities improve significantly, and in some cases, some LLMs are also capable of handling various types of information (e.g., images or audio) (Liu et al., 2023a; Chen et al., 2023; Wang et al., 2023d; Lyu et al., 2023; Li et al., 2023b). These improvements render some external tools that may not be necessary for some LLMs in the future. Therefore, we have annotated the reasons why LLMs need to use these tools to solve user problems.
@@ Appendix B Test Set for Evaluation on the Awareness of Tool Usage
To assess to what extent LLMs are aware of their limitations, we construct the dataset for the awareness of tool usage by merging the positive samples from a subset of ToolE dataset and the negative samples from some subsets of public datasets. To exclude the ambiguous situation where the queries in ToolE can be solved either with or without the assistance of external tools, we manually check the output of the queries by feeding them directly into the LLMs and retain those whose responses are unsatisfactory(i.e. the model apologizes, the response contains errors, and etc). In this way, we verify that solving the queries in our selected subset(containing 515 samples) is beyond the capacity of the existing LLMs and the queries can therefore be treated as positive samples when evaluating the awareness of tool usage. Meanwhile, we collect negative samples of equal size from the public datasets of daily conversations (Wang et al., 2022; Zhou et al., 2023) and common sense (Talmor et al., 2019), the queries of which are expected to be answered by LLMs without the help of external tools.
@@ Appendix E Failure Case Study
6 |
Converting 85 F to Celsius. |
Yes. Converting Fahrenheit to Celsius requires a conversion formula. Although I possess knowledge of this calculation, I do not have the capability to perform mathematical operations. I would need to use an external tool, such as a calculator or programming language, to accurately convert the temperature. |

### 23 Feb 2024 v5

# MetaTool Benchmark for Large Language Models: Deciding Whether to Use Tools and Which to Use

@@ 1 Introduction
Tool-empowered large language models (LLMs) (Qin et al., 2023a; b; Patil et al., 2023; Ruan et al., 2023; Cao et al., 2023; Zhou et al., 2023a) have recently attracted widespread attention. An important milestone for LLMs marching toward intelligent agents (Park et al., 2023; Li et al., 2023e) is the flexible use of tools (e.g., APIs (Qin et al., 2023b; Rapid, 2023) and plugins (OpenAI, 2023d)) to fulfill users’ requirements. By utilizing tools, LLMs can obtain real-time data, such as getting the latest weather forecast (GPTStore, 2023); enhance interactions with users, like helping users book flight tickets (Deng et al., 2023); and better deal with uncertain questions by querying knowledge bases (Li et al., 2023c; Hu et al., 2023) or Internet (Lazaridou et al., 2022). Moreover, LLMs can also leverage specific tools to process multimodal information, thereby acquiring the same capabilities as multimodal models (Zhang et al., 2023; Yan et al., 2023; Yuan et al., 2023b). The capacity to use tools enables LLMs to break through their own limitations, acquire external information, and thereby make more accurate and effective responses, providing users with better service.
Previous research has focused on how to enhance the ability of LLMs to use tools, including training models with instruction related to tool usage (Qin et al., 2023b; Tang et al., 2023; Schick et al., 2023), or augmenting the model’s problem-solving capabilities for domain-specific tasks through external APIs (Yang et al., 2023c). A typical process of employing LLMs to use tools is illustrated in Figure 1. Initially, users input a question (i.e., query) that triggers the tool usage. Based on prior research (Yang et al., 2023c; Qin et al., 2023b), under the ReAct (Yao et al., 2022) prompt approach, the process of using tools can be divided into four stages: Firstly, LLMs consider whether to employ a tool (①) and if so, which tools to select (②). The tool selection process involves directly having LLMs choose from a provided tool list (Yang et al., 2023c) or selecting via a retriever (Qin et al., 2023b). Next, LLMs configure the users’ input as tool parameters (③), then handle the results from the tool (④), and finally return the outcomes to the user.
@@ 2.1 Preliminary & Required Abilities
In this section, we first introduced the composition of the ToolE dataset, outlining how we generated user queries related to tools. Subsequently, we explained how we set up the evaluation tasks, including tool usage awareness and tool selection. The evaluation mainly requires LLMs with the following properties and abilities: (1) Less hallucination and sycophancy. The awareness of tool usage can reflect the truthfulness about whether an LLM has a clear understanding of its capabilities (e.g., realizing its capability limitation about what problems it cannot solve well and using tools for assistance), thereby helping to mitigate issues of hallucination (Ji et al., 2023; Sun et al., 2024) and sycophancy (Wei et al., 2023). (2) Recommendation and retrieval. Moreover, existing research has tentatively explored the potential of LLMs in applications like LLM-based recommendation systems (e.g., tool recommendation for users) (Gao et al., 2023; Wang et al., 2023e; Dai et al., 2023). In LLM-as-agent scenarios, LLMs usually need to select the specific tool according to the text description (Park et al., 2023; Shen et al., 2023; Ruan et al., 2023), actually is a kind of information retrieval (Sun et al., 2023), making the ability of tool selection crucial. (3) Task-level abilities. In MetaTool, we set four tasks as shown in Table 12. Incorporating similar tools for selection (i.e., Task 1) requires a high-level semantic comprehension for LLMs, and tool selection in specific scenarios tests the flexibility of LLMs when using tools in different scenarios (e.g., finance (Wu et al., 2023a) and biomedical domain (Zhang et al., 2023; Wang et al., 2023c)). Task 3 aims to explore the internal hallucination and reliability extent of LLMs when using tools and Task 4 is designed to evaluate the inference ability (e.g., order of using multiple tools) (Creswell et al., 2022) of LLMs.
@@ 2.2.1 Dataset Generation
Multi-tool queries generation. Unlike single-tool queries, we generate multi-tool queries after addressing the overlapped issue because it is challenging to map the original labels to new labels in multi-label (i.e., multi-tool) situations. Here, we only consider queries related to two tools. We observe that if we obtain combinations of two tools by iterating through all the tools (i.e., C iterations, where is the size of the toolset), there would be many tool combinations that are not practical (i.e., rarely encountered in daily life, such as the combination of fortune-telling tools and currency exchange tools). Therefore, we select the top 15 most popular tools from the toolset, and for each pair of tools, we generate 5 queries. We determine the popularity of a tool based on the number of tools it is merged with, as shown in Appendix A.5. The multi-tool queries we generate can be divided into two types: The first category pertains to situations where tools are employed in parallel, indicating that the utilization of each tool operates independently of the others. The second category deals with cases where tools are used causally, signifying that the deployment of one tool may be contingent upon the outcomes of a preceding tool. Detailed prompt templates can be found in Appendix D.2. Similar to single-tool queries, we also manually verified multi-tool queries to ensure the combination of tools was reasonable and the query of the tool corresponded to the tool description.
@@ 2.3.1 Awareness of Tool Usage
In this part (i.e., Thought (①)), we aim to investigate the awareness of tool usage of LLMs; that is, whether LLMs can resort to external tools when they encounter problems they cannot solve. To this end, we need to construct the test set with both positive and negative samples. Positive samples are the queries that can not be solved by LLMs themselves and need tool usage, whereas negative samples are queries that can be directly solved by LLMs and therefore do not necessitate tool usage. For positive samples, we selected a subset of samples from ToolE and conducted manual validation to confirm whether they would trigger LLMs to use the tool (the process of which is detailed in the Appendix B). As for negative samples, we select three recent instruction datasets, including instructions about downstream tasks (Wang et al., 2022), common-sense questions (Talmor et al., 2019), and high-quality instructions used in LIMA (Zhou et al., 2023b). Similarly, we conducted manual verification to ensure that these requests can be resolved by LLMs’ intrinsic capabilities. Specifically, we use the prompt with a query to inquire the LLMs whether need to employ a tool or not, and the output of LLMs should be either "yes" or "no".
@@ 2.3.2 Tool Selection
Sub-task 2: tool selection in specific scenarios. The objective of this task is to simulate how LLMs perform using tools when they act as controllers of a system (Shen et al., 2023) faced with different scenarios. As LLMs are widely applied across various domains like biomedical domain (biomedgpt) and educational domain (Kasneci et al., 2023), in scenarios where the system caters to diverse demographics or professions (e.g., software engineer (Qian et al., 2023)), its set of tools also varies. This task allows us to explore the performance disparities of LLMs in selecting different kinds of tools, essentially highlighting a form of bias inherent to LLMs (Ferrara, 2023). In such cases, this task examines how effectively LLMs utilize the tools. Given a query with its label , we specify the tool list containing candidates according to its corresponding scenarios. This task consists of two types of scenarios: the first one is the popularity-related scenario, and the second one is the group-related scenario. For the popularity-related scenario, we have selected the 5, 10, and 15 most popular tools based on the number of tools it’s merged with (refer to Table 9 in the Appendix A.5 for details.) to construct the tool list. As for the group-related scenario, we chose six usual occupations or identities and manually curated a tool list consisting of the 10 most relevant tools for each of them (see Table 10 in the Appendix A.5 for details).
Sub-task 3: tool selection with possible reliability issues. The reliability of LLMs’ tool selection is of utmost importance. However, issues like hallucination (Ji et al., 2023) and sycophancy (Wei et al., 2023) within LLMs’ responses that will negatively impact their selection of tools. Therefore, we introduce sub-task 3. In this task, given a query and its corresponding tool , we need to construct the tool list and ensure . This aims to assess whether LLMs can answer questions honestly and avoid issues like choosing non-existent tools or selecting unrelated tools. It should be noted that this task closely resembles real-world scenarios, as not all existing tools capable of addressing user queries are present in the tool list controlled by LLMs. To be specific, we obtain the embedding of ’s description and get the top-k similar tools about as the way in Task 1. Then we randomly sample tools from the rest tool set to construct , denoted as where and . Overall, we remove the ground-truth tool of query and s most similar tools to keep the tools in not related to as much as possible.
@@ A.3 Guidelines for human validation
Calculation-related tools. ToolE contained numerous tools related to calculations. For simple calculations (e.g., ‘What is the value of sin 30 degrees?’ or ‘7 * 9 = ?’), LLMs can perform them without the need for a tool. However, for complex calculations, recent research (Yuan et al., 2023a) suggests that LLMs still perform poorly. For queries corresponding to calculation-related tools, we removed queries involving simple calculations and retained those involving complex calculations.
@@ A.4 Flexibility of ToolE
Sometimes, when an LLM undergoes specialized training in a particular domain (Wu et al., 2023b; Wang et al., 2023a; Qi et al., 2023; Wang et al., 2023b; Yang et al., 2023a), its capabilities improve significantly, and in some cases, some LLMs are also capable of handling various types of information (e.g., images or audio) (Liu et al., 2023; Chen et al., 2023; Wang et al., 2023d; Lyu et al., 2023; Li et al., 2023b). These improvements render some external tools that may not be necessary for some LLMs in the future. Therefore, we have annotated the reasons why LLMs need to use these tools to solve user problems.
@@ Appendix B Test Set for Evaluation on the Awareness of Tool Usage
To assess to what extent LLMs are aware of their limitations, we construct the dataset for the awareness of tool usage by merging the positive samples from a subset of ToolE dataset and the negative samples from some subsets of public datasets. To exclude the ambiguous situation where the queries in ToolE can be solved either with or without the assistance of external tools, we manually check the output of the queries by feeding them directly into the LLMs and retain those whose responses are unsatisfactory(i.e. the model apologizes, the response contains errors, and etc). In this way, we verify that solving the queries in our selected subset(containing 515 samples) is beyond the capacity of the existing LLMs and the queries can therefore be treated as positive samples when evaluating the awareness of tool usage. Meanwhile, we collect negative samples of equal size from the public datasets of daily conversations (Wang et al., 2022; Zhou et al., 2023b) and common sense (Talmor et al., 2019), the queries of which are expected to be answered by LLMs without the help of external tools.